OpenStudy (anonymous):

Evaluate the definite integral. The integral from 0 to a of x*sqrt(a^2-x^2)dx

OpenStudy (anonymous):

\[\int\limits_{0}^{a}x \sqrt{a^2-x^2}dx\]

OpenStudy (tkhunny):

Have you considered a substitution?

OpenStudy (anonymous):

Yeah. I set u=a^2-x^2

OpenStudy (primeralph):

Yeah, that can work. Or trig.

OpenStudy (tkhunny):

Okay. Next? du = ??

OpenStudy (anonymous):

I don't know trig subsitution.

OpenStudy (anonymous):

du=2a-2xdu

OpenStudy (anonymous):

Or would a^2 just be left out because it is a constant?

OpenStudy (tkhunny):

Close. a is a constant. That should go away. Why is there a "du" on the right side? Give it another go.

OpenStudy (anonymous):

du=-2xdx then you would divde by -2 and get -1/2du=xdx

OpenStudy (tkhunny):

Okay. Next...

OpenStudy (anonymous):

=\[=\int\limits_{0}^{a}\frac{ -1 }{ 2 }(\sqrt{u})du\]

OpenStudy (tkhunny):

No good. The limits are still x-limits. You have two options: 1) Change the limits to limits of "u", or 2) Solve the indefinite integral and change the variable back to x before evaluating the definite integral.

OpenStudy (anonymous):

I solved the indefinite integral and changed it back to x. I got \[\frac{ 1 }{ 2 }(\frac{ 2 }{ 3 }(a^2-x^2)^{\frac{ 3 }{ 2 }}+C)\]

OpenStudy (anonymous):

Then wouldn't you just evaluate it at a and 0? then subtract where you evaluated it at 0 from where you evaluated it at a?

OpenStudy (tkhunny):

That seems mostly okay. Simplify the expression and figure out where the negative sign went.

OpenStudy (anonymous):

Oh, I just forgot to type it. It's still there in my work.

OpenStudy (anonymous):

I got \[\frac{ -1 }{ 3 }a^\frac{ 4 }{ 3 }\] as my final answer.

OpenStudy (tkhunny):

It should not be negative. How did the 3 int he exponent get into the denominator of the exponent? Try again. The first term is zero. You should get only the second term.

OpenStudy (tkhunny):

Reality check. Look at the original argument. For positive 'a', can you find anything negative about that expression? It better not be a negative result.

OpenStudy (anonymous):

so would it be \[\frac{ 1 }{ 3 }a^3\]

OpenStudy (tkhunny):

Why is it NOT \(\dfrac{1}{3}|a|^{3}\)?